# Strong Screening Rules for Group-based SLOPE Models

### v1

@@ 5.2 Real data experiments
Both gSLOPE and SGS were fitted with their FDR-control parameters set to $0.01$ and for SGS $\alpha=0.99$. Each model was fitted along a path of $100$ regularization parameters, with $\lambda_{1}$ set as described in Appendices A.3 and B.4, and $\lambda_{100}=0.01\lambda_{1}$. The data was $\ell_{2}$ standardized and no intercept was used.
Results.
For both gSLOPE and SGS, the screening rules lead to considerably faster runtimes (Figure 5). In particular, the use of the screening rules significantly benefits the gSLOPE approach. With the reduced input feature space, any issues with convergence of the fitting algorithm are substantially decreased. For the cancer dataset, $186$ model fits (out of the $900$) did not reach convergence with no screening, compared to only $16$ with screening. A similar trend occurred for the colitis dataset. As gSLOPE applies no variable penalization, it is forced to fit all variables within a group. For datasets with large groups, such as those considered here, this leads to a problematic fitting process which can include many noisy variables.
@@ Results.
The analysis of the real data further illustrates the benefits of the bi-level screening to the runtime and performance of SGS. Figure 1(b) illustrates that for these large genetics datasets, the bi-level screening allows the input dimensionality for SGS to be reduced to a much greater extent than by just group screening.
Both screening rules can drastically reduce the input dimensionality as seen in Table 2. For the colitis dataset, the SGS screening rule reduces the input dimensionality to just $2\%$ of the full space. This comes without the cost of solution consistency. The estimated model coefficients with and without screening were very close to each other, with their $\ell_{2}$ distances of order $10^{-8}$ (Appendices F.2 and F.4).
@@ 5.3 KKT violations
The screening rules rely on assumptions that can fail. In these cases, KKT checks are implemented to ensure that no active variables are excluded from $\mathcal{E}_{v}$. A KKT violation occurs when a variable is not included in the screening set, $\mathcal{S}_{v}$, but should be according to the KKT optimality conditions (Equation 2), in which case it is added into the fitting set, $\mathcal{E}_{v}$. KKT violations are very rare for gSLOPE (Figure 6), occurring on the simulated data infrequently toward the start of the path. On the real data, no KKT violations occurred for gSLOPE.
(a) Group violations for gSLOPE and variable violations for SGS, fitted to linear models, averaged over all cases of $p$ and the correlation ($\rho$).
(b) Variable violations for SGS applied to the cancer and colitis datasets, averaged over the nine pathway collections.
Figure 6: The proportion of KKT violations relative to the full input space, as a function of the regularization path.
For SGS, KKT violations are more common, as the additional layer of screening carries with it additional assumptions. Additionally, the subdifferential term in Equation 7 was chosen to be as small as possible, leading to tighter screened sets. The increased violations are reflective of this. On both the simulated and real datasets, SGS recorded an increasing number of KKT violations as the fitted models became denser (Figure 6). The shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path. The trend of increasing KKT violations with denser models is also observed for the SLOPE strong screening rule [17].
Despite the increased number of KKT violations for SGS, which amounts to the additional computational cost of performing the KKT checks and repeated sub-fitting in Algorithm 1, the overall screening procedure still results in substantial runtime improvements, as evidenced in Tables 1 and 2. This shows that screening rules can provide improvements to computational cost, even with assumptions that are often violated.
@@ 6 Discussion
In this manuscript, we have developed strong screening rules for group-based SLOPE models: Group SLOPE and Sparse-group SLOPE, neither of which have any previous screening rules. The screening rule for gSLOPE screens out irrelevant groups before fitting. The screening rules for SGS perform bi-level screening, based on our proposed sparse-group screening framework.
The two proposed screening rules differ from the existing SLOPE screening rules both in construction and in final outcome. gSLOPE only performs screening through the groups, forcing it to keep all variables within the screened groups, even noisy ones, in the optimization process. SGS, similarly to SLOPE and in contrast to gSLOPE, performs variable screening after its group screening step. As additional assumptions are required for SGS to perform bi-level selection, its computational efficiency is not as effective as the one observed with the SLOPE screening rules [17]. The superiority of SGS in variable selection and predictive performance when group information is available overcomes the additional computational burden [14].
@@
Through comprehensive analysis of synthetic and real data, we illustrate that the screening rules lead to dramatic improvements in the runtime of gSLOPE and SGS models, as well as for group-based OSCAR models (Appendix E). This is achieved without affecting model accuracy. This is particularly important in datasets where $p\gg n$, such as genetics ones, which is the main motivation behind SLOPE [4]. The screening rules presented in this manuscript allow group-based SLOPE, and by extension, group-based OWL models, to achieve computational fitting times more in line with lasso-based models, making them more widely accessible.
Limitations.
@@
One of the limitations of our proposed screening rules for SGS is the number of assumptions made. A future direction includes the exploration of alternative strong screening rules that require fewer assumptions. Alternatively, the development of safe rules, which guarantee that only inactive variables are discarded, could be an avenue for future exploration, potentially incorporating safe and strong rules together in a hybrid rule [32, 37]. A comparison between safe and the proposed strong rules developed in this manuscript could provide further insight into which type of screening is most beneficial for non-separable penalties.
Appendix
[…]
@@ A.1 Penalty weights
The penalty weights for gSLOPE were derived to provide group FDR-control under orthogonal designs [6]. For the FDR-control parameter $q_{g}\in(0,1)$, they are given by (where the indexing corresponds to the sorted groups)
@@
where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom. A relaxation to this sequence is applied in [6], to give
@@
A.2 Theory
@@
The proof is similar to that of Theorem 2.7 in [6], where the subdifferential of gSLOPE is derived under equal groups. It is derived here under more general terms. The subdifferential needs to be derived under two cases:
@@
Case 1: For inactive groups, we consider the subdifferential at zero. The subdifferential of a norm at zero is given by the dual norm of the unit ball [27],
@@
The dual norm for gSLOPE is given by [6]
@@
where $\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]=\{x\in\mathbb{R}^{m}:\text{cumsum}(|x|_{\downarrow}-w)\preceq\mathbf{0}\}$ is the unit ball of the dual norm to $J_{\text{slope}}$ [4]. Using this, the subdifferential at zero for the inactive groups, $\mathcal{Z}$, is given by
@@
Since the set $\{d^{(g)}:d\in D_{g}\}$ is open in $\mathbb{R}^{p_{g}}$ and contains zero, by Corollary G.1 in [6], it follows that $x^{(g)}\in\partial f_{g}(b^{(g)})$ for $f_{g}:\mathbb{R}^{p_{g}}\rightarrow\mathbb{R},f_{g}(x)=w_{g}\sqrt{p_{g}}\|x\|_{2}$. Now, for $g\leq s$, $f_{g}$ is differentiable in $\beta^{(g)}$, giving
@@
Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then, plugging in $h(\lambda_{k+1})=([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}$ gives
@@
Since $\text{cumsum}(y)\succeq\text{cumsum}(x)\iff y\succeq x$ [17], we only need to show for a group $g$,
@@
A.3 Path start derivation
@@
$\lambda\succeq\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w),$
[…]
@@
This can be verified by noting that $\lambda_{1}=J^{*}_{\text{gslope}}(\nabla f(\mathbf{0});w)$ [22]. Now, $J^{*}_{\text{gslope}}(x;w)=J^{*}_{\text{slope}}([x]_{\mathcal{G},-0.5};w)$ [6]. The dual norm of SLOPE is given by [23]
@@
Therefore, $\lambda_{1}$ is as before.

### v2

@@ 6.2 Real Data Experiments
Both gSLOPE and SGS were fit with their FDR-control parameters set to $0.01$ and for SGS $\alpha=0.99$. Each model was applied along a path of $100$ regularization parameters, with $\lambda_{100}=0.01\lambda_{1}$. Table A1 describes the algorithmic parameters used for ATOS.
Results
For both gSLOPE and SGS, the screening rules led to considerably faster runtimes for all datasets (Figure 5). The screening was more effective for SGS under a linear model and for gSLOPE under a logistic model. The active set sizes for gSLOPE tended to be smaller for the logistic models, allowing for larger feature reduction (Table 2). For SGS, this trend went the other way; the active sets were smaller for the continuous datasets. Another explanation for the reduced screening efficiency for gSLOPE under linear models was due to the grouping structure. For the continuous responses, the groups were generated using K-means clustering, leading to fewer groups and less opportunity for dimensionality reduction (Table A8). With fewer groups, gSLOPE is less likely to discard full groups, as they may contain some signal.
The feature reduction provided by our screening rules led to the alleviation of convergence issues (Table F.5). SGS failed to converge for three datasets without screening and none with screening. gSLOPE experienced failed convergences across all datasets without screening, while with screening, it only failed to converge for three datasets, each showing fewer instances of failure. As gSLOPE applies no variable penalization, it is forced to fit all variables within a group. For datasets with large groups, such as those considered here, this leads to a problematic fitting process which can include many noisy variables. Our screening rules help gSLOPE partially overcome this issue, leading to large computational savings and better solution optimality.
@@ Results
The analysis of the real data further illustrates the benefits of the bi-level screening to the runtime and performance of SGS (for synthetic data, see Figure 1). Figure 6 illustrates that for the cancer and celiac datasets, the bi-level screening allows the input dimensionality for SGS to be reduced to a much greater extent than by just group screening (see Figure A17 for the other datasets).
Both screening rules drastically reduce the input dimensionality on all the real datasets (Table 2). On average, for the logistic real data models, the SGS screening rules reduced the input to just $7\%$ of the total space. This comes without affecting solution consistency (Appendices F.2 and A16).
Figure 5: The ratio of no screen time to screen time ($\uparrow$) of gSLOPE and SGS applied to the real datasets, for fitting $100$ path models, split into response type. The horizontal grey line represents no screening improvement.
Table 2: The cardinality of the active ($\mathcal{A}$) and fitting ($\mathcal{E}$) sets for gSLOPE and SGS, averaged across all synthetic and real data cases, split into model type. For gSLOPE, the cardinality is for the group sets, and for SGS the variable ones. Dim. is the input dimensionality. For synthetic, it was $p=2750$ and $m=210$.
Synthetic Real
[…]
gSLOPE Lin. $55\scriptstyle{\pm 1}$ $76\scriptstyle{\pm 1}$ $112\scriptstyle{\pm 3}$ $168\scriptstyle{\pm 3}$ $240$
gSLOPE Log. $71\scriptstyle{\pm 1}$ $97\scriptstyle{\pm 1}$ $49\scriptstyle{\pm 2}$ $83\scriptstyle{\pm 3}$ $1634$
SGS Lin. $178\scriptstyle{\pm 3}$ $364\scriptstyle{\pm 6}$ $378\scriptstyle{\pm 37}$ $1139\scriptstyle{\pm 87}$ $14488$
SGS Log. $230\scriptstyle{\pm 3}$ $472\scriptstyle{\pm 6}$ $526\scriptstyle{\pm 5}$ $979\scriptstyle{\pm 13}$ $13734$
For additional metrics and comparison to the SLOPE strong rule for the real datasets, see Figures A14, A15, and A16. Our rules for gSLOPE and SGS are found to offer similar levels of runtime improvements and feature reduction to that of SLOPE (Larsson et al., 2020).
Figure 6: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for group-only and bi-level screening applied to SGS, plotted along the regularization path for the cancer and celiac datasets.
@@ 6.3 KKT Violations
As with any strong screening rule, our approach depends on assumptions that may fail. When this happens, KKT checks are used to ensure no active variables are excluded, with violations added to $\mathcal{E}_{v}$. For SGS, KKT violations occur at a variable-level (Appendix B.4) and for gSLOPE at a group-level (Appendix A.4).
KKT violations are very rare for gSLOPE (Figure 7), occurring on the simulated data infrequently toward the start of the path. On the real data, only a single dataset had violations, and the number of violations was very small (Table F.5).
For SGS, KKT violations are more common, but still infrequent (Figure 7 and Tables A9) due to additional assumptions in the second layer of screening. In Equation 7, minimizing the subdifferential term leads to tighter screened sets, contributing to these violations. In Figure 7, the number of violations increases as a function of the model density for SGS, mirroring the log-linear shape of the regularization path. This pattern is also seen in the strong rule for SLOPE (Larsson et al., 2020).
Despite these violations and the computational cost of the checks and refitting (Algorithm 1), the overall screening process yields significant runtime improvements in all cases (Table 1 and Figure 5).
Figure 7: The proportion of KKT violations relative to the full input for gSLOPE and SGS, under linear models, averaged over all synthetic data cases. This is shown as a function of the regularization path.
@@ 7 DISCUSSION
In this manuscript, we have developed strong screening rules for group-based SLOPE models using our new sparse-group strong screening framework: Group SLOPE and Sparse-group SLOPE, neither of which have any previous screening rules. Our proposed screening rules differ from the existing SLOPE strong rule both in construction and in outcome. The screening rule for gSLOPE screens out irrelevant groups before fitting. The screening rules for SGS perform bi-level screening. Our rules apply to the wider class of OWL models, including group-based OSCAR models.
SLOPE models are finding increasing use in genetics and machine learning, with SGS found to have superior disease prediction performance over other penalized methods (Feser and Evangelou, 2023). Our screening rules will make the group-based versions more accessible by reducing their computational burden. This will allow practitioners to utilize their FDR properties, facilitating their widespread use across various fields.
@@
Through comprehensive analysis of synthetic and real data, we illustrate that the screening rules lead to dramatic improvements in the runtime of gSLOPE and SGS models, as well as for group-based OSCAR models (Appendix E). This is achieved without affecting model accuracy. This is particularly important in datasets where $p\gg n$, such as genetics ones, which is the main motivation behind the proposal of SLOPE (Bogdan et al., 2015). The screening rules presented in this manuscript allow group-based SLOPE, and more generally group-based OWL models, to achieve computational fitting times more in line with their lasso-based counterparts. The screening rules also helped gSLOPE and SGS overcome convergence issues in large datasets, improving solution optimality.
In our data studies, we have not discovered any scenario where our screening rules did not perform better than no screening. In each case, the rules greatly reduce the input dimensionality and speed up the computational runtime.
Limitations
Our screening rules, as any strong rule, rely on assumptions. For both gSLOPE and SGS, Lipschitz assumptions were used that are consistent with those used in the strong screening framework (Tibshirani et al., 2010). For gSLOPE, a Lipschitz assumption was used to derive Proposition 3.3, while for SGS, a separate Lipschitz assumption was made for each layer of screening (Propositions 4.2 and 4.3).
Violations of these assumptions are checked for using the KKT conditions. The SGS KKT checks (Appendix B.4) led to an increased number of violations, as the checks were overly conservative. However, the overall amount of KKT violations for both models was still relatively small, suggesting that these assumptions are not overly restrictive.
An attempt was made to derive less conservative checks (Appendix B.4.3), performed directly on the variables without an initial group check. These were found to be too lenient, incorrectly missing violations. Future work can consider alternative KKT checks.
@@
An additional future direction includes the development of safe rules, which guarantee that only inactive variables are discarded. These could be incorporated into a hybrid scheme together with our strong rules (Zeng et al., 2021; Wang and Breheny, 2022). Deriving safe rules would facilitate a comparison between them and our proposed strong rules, offering further insight into which type of screening is most effective for both non-separable and sparse-group norms.
Acknowledgements
We would like to thank the anonymous reviewers for their valuable comments. This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) through the Modern Statistics and Statistical Machine Learning (StatML) CDT programme, grant no. EP/S023151/1.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes, we have defined our approach and the problem setting.]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes, we have described the complexity of the fitting algorithm.]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes. We have presented our assumptions in the results and have discussed the limitations of these in the discussion.]
    2. (b)
      Complete proofs of all theoretical results. [Yes. Proofs have been provided in the Appendix for all of our results.]
    3. (c)
      Clear explanations of any assumptions. [Yes, we have extensively discussed our assumptions and any corresponding limitations.]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes, code has been provided to reproduce any figure.]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes, we have given the full simulation set up in Table A1.]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes, we have described our statistical measures and our error bars.]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes, we have cited our data and model sources.]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Yes, this is included in the code supplementary materials.]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Strong Screening Rules for Group-based SLOPE Models:
Supplementary Materials
@@ A.1 Penalty Weights
The penalty weights for gSLOPE were derived to provide group FDR-control under orthogonal designs (Brzyski et al., 2019). For the FDR-control parameter $q_{g}\in(0,1)$, they are given by (where the indexing corresponds to the sorted groups)
@@
where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom. A relaxtion to this sequence is applied in Brzyski et al. (2019) to give
@@
A.2 SLOPE Subdifferential
The subdifferential for SLOPE was derived in Larsson et al. (2020) and as it is a vital part of our arguments, it is reproduced here for ease of reference.
Define the function $R:\mathbb{R}^{p}\rightarrow\mathbb{N}^{p}$ that returns the ranks of the absolute values of its input and the set $\mathcal{C}_{i}(\beta)=\{j\in\{1,\ldots,p\}\;:\;|\beta_{i}|=|\beta_{j}|\}$. Then, the subdifferential is given by (Larsson et al., 2020)
$\partial J_{\text{slope}}(\beta;v)=\begin{cases}\left\{x\in\mathbb{R}^{|\mathcal{C}_{i}|}:\text{cumsum}(|x|_{\downarrow}-v_{R(x)_{\mathcal{C}_{i}}})\preceq 0\right\}&\text{if}\;\beta_{\mathcal{C}_{i}}=0,\\ \Big{\{}x\in\mathbb{R}^{|\mathcal{C}_{i}|}:\text{cumsum}(|x|_{\downarrow}-v_{R(x)_{\mathcal{C}_{i}}})\preceq 0\\ \;\;\;\;\;\;\text{and}\;\;\sum_{j\in\mathcal{C}_{i}}(|x_{j}|-v_{R(s)_{j}})=0\\ \;\;\;\;\;\;\text{and}\;\;\text{sign}(\beta_{\mathcal{C}_{i}})=\text{sign}(x)\Big{\}}&\text{otherwise.}\end{cases}$
The primary use of the subdifferential in this manuscript is the zero condition (Equation 5).
A.3 Theory
@@
The proof is similar to that of Theorem 2.7 in Brzyski et al. (2019), where the subdifferential of gSLOPE is derived under equal groups. It is derived here under more general terms. The subdifferential needs to be derived under two cases:
@@
Case 1: For inactive groups, we consider the subdifferential at zero. The subdifferential of a norm at zero is given by the dual norm of the unit ball (Schneider and Tardivel, 2022),
@@
The dual norm for gSLOPE is given by (Brzyski et al., 2019)
@@
where $\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]=\{x\in\mathbb{R}^{m}:\text{cumsum}(|x|_{\downarrow}-w)\preceq\mathbf{0}\}$ is the unit ball of the dual norm to $J_{\text{slope}}$ (Bogdan et al., 2015). Using this, the subdifferential at zero for the inactive groups, $\mathcal{Z}$, is given by
@@
Since the set $\{d^{(g)}:d\in D_{g}\}$ is open in $\mathbb{R}^{p_{g}}$ and contains zero, by Corollary G.1 in Brzyski et al. (2019), it follows that $x^{(g)}\in\partial f_{g}(b^{(g)})$ for $f_{g}:\mathbb{R}^{p_{g}}\rightarrow\mathbb{R},f_{g}(x)=w_{g}\sqrt{p_{g}}\|x\|_{2}$. Now, for $g\leq s$, $f_{g}$ is differentiable in $\beta^{(g)}$, giving
@@
Suppose we have $\mathcal{B}\neq\varnothing$ after running the algorithm. Then, plugging in $h(\lambda_{k+1})=([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}$ gives
@@
Since $\text{cumsum}(y)\succeq\text{cumsum}(x)\iff y\succeq x$ (Larsson et al., 2020), we only need to show for a group $g$,
@@
A.4 KKT Checks
To check whether a group has been correctly discarded during the screening step, the KKT conditions for gSLOPE are checked. They are given by
$\mathbf{0}$ $\in\nabla f(\beta)+\lambda\partial J_{\text{gslope}}(\beta;w)$
$\implies-\nabla f(\beta)$ $\in\lambda\partial J_{\text{gslope}}(\beta;w).$
Hence, we are checking whether the gradient of the loss function sits within the set of the gradient of the penalty. As we are only interested in identifying incorrectly discarded groups, we require only to check the subdifferential condition at zero. Hence, a violation occurs if a group is discarded but
$-\nabla f(\beta)\notin\lambda\partial J_{\text{gslope}}(\mathbf{0};w_{\mathcal{G}_{\mathcal{Z}}})$
$\implies$ $-\nabla f(\beta)\notin\left\{x\in\mathbb{R}^{\operatorname{card}\mathcal{G}_{\mathcal{Z}}}:[x]_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;\lambda w_{\mathcal{G}_{\mathcal{Z}}})\right\}$
$\implies$ $[\nabla f(\beta)]_{\mathcal{G}_{\mathcal{Z}},-0.5}\notin\partial J_{\text{slope}}(0;\lambda w_{\mathcal{G}_{\mathcal{Z}}})$
$\implies$ $\text{cumsum}(([\nabla f(\beta)]_{\mathcal{G}_{\mathcal{Z}},-0.5})_{\downarrow}-\lambda w_{\mathcal{G}_{\mathcal{Z}}})\succ 0.$
A.5 Path Start Proof
Proof of Proposition 3.4.
@@ A.5 Path Start Proof
$\lambda\succeq\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w).$
Picking the maximum possible $\lambda$ such that this holds yields
@@
This can be verified by noting that $\lambda_{1}=J^{*}_{\text{gslope}}(\nabla f(\mathbf{0});w)$ (Ndiaye et al., 2016b). Now, $J^{*}_{\text{gslope}}(x;w)=J^{*}_{\text{slope}}([x]_{\mathcal{G},-0.5};w)$ (Brzyski et al., 2019). The dual norm of SLOPE is given by (Negrinho and Martins, 2014)
@@
Therefore, $\lambda_{1}$ is as before. ∎